# A Neural Network Enhanced Born Approximation for Inverse Scattering

### v2

@@ Abstract
Time-harmonic, acoustic inverse scattering concerns the ill-posed and nonlinear problem of determining the refractive index of an inaccessible, penetrable scatterer based on far field wave scattering data. When the scattering is weak, the Born approximation provides a linearized model for recovering the shape and material properties of a scatterer. We develop two neural network algorithms–Born-CNN (BCNN) and CNN-Born (CNNB)–to correct the Born approximation when the scattering is not weak. BCNN applies a post-correction to the Born reconstruction, while CNNB pre-corrects the data. Both methods leverage the Born approximation’s excellent fidelity in weak scattering, while extending its applicability beyond its theoretical limits. CNNB particularly exhibits a strong generalization to noisy and absorbing scatterers. Based on numerical tests, our approach provides alternative data-driven methods for obtaining the refractive index, extending the utility of the Born approximation to regimes where the traditional method fails.
@@ 1 Introduction
We shall study a particular inverse scattering problem: the inverse medium problem for the Helmholtz equation. In this case, it is desired to reconstruct the refractive index of a bounded penetrable scatterer from far field acoustic data. The major complication is that this inverse scattering problem is both ill-posed and nonlinear [6, Theorem 4.21, page 448]. Current approaches can be divided into two broad classes: 1) quantitative methods and 2) qualitative methods. A quantitative method attempts to reconstruct the scatterer directly. Often this involves a non-linear and regularized optimization problem that seeks a reconstruction corresponding to a far field pattern that matches measurement data. This is computationally intensive and can fail due to local minima [6]. Another quantitative approach is to assume that the scattering is weak and to use the inverse Born approximation. We shall discuss this in more detail shortly since correcting this approach is the subject of our paper. In contrast to these quantitative methods, the goal of a qualitative method such as the Linear Sampling Method [3, 6] is to approximate the support or boundary of the scatterer. Approaches of this type do not involve optimization or the solution of the forward problem. They do not require strong a priori assumptions, but cannot directly distinguish different materials in the field of view.
@@
The forward problem that maps a known refractive index to the predicted far field pattern is given by a Neumann series in a certain integral operator, provided an appropriate norm of the integral operator is less than one. Conditions for this to occur have been derived in several cases (see, for example, [11] in the seismic context) and we recall a simple sufficient condition in Section 2. Selecting only the first term in this series defines a linear map, the Born approximation, from the contrast to the far field pattern.
@@
Turning to the inverse problem, the inverse Born approximation can be understood as inverting the linearized Born approximation. Inverting this linear map is ill-posed, but removes the difficulty of nonlinearity and provides an avenue to solve the inverse problem approximately using regularization to restore stability. An early reference for this technique is  [2], while applications and computational techniques are described in [9, 8]. The inverse Born approximation has been widely applied to neutron scattering, medical imaging, and seismic inverse problems [11]. Within the weak scattering approximation, there has been a great deal of work to incorporate higher-order expansions into the forward and inverse Born approximation (see the review of Moskow and Schotland [20]). Our goal in this paper is to correct the Born approximation and extend its applicability beyond the weak scattering limit using neural networks as correctors.
@@
Another machine learning approach to correcting the Born approximation is the statistical approximation error correction used in [13]. That paper shows that the Born approximation can be extended outside the weak scattering approximation by a suitable training approach if the scatterer is well represented in the training data. The approach of [13] is based on Bayesian statistics and not on neural networks.
Viewing the map from the far field to the regularized reconstruction of the scatterer as an unknown nonlinear function, we can appeal to the universal approximation property (first proved in [7]) of neural networks to approximate this map. This is the approach in [10] where a novel network is trained to solve the inverse scattering problem. Numerical results show excellent reconstructions when the Born approximation holds.
There have been several other developments in designing specialized neural network-based solvers for the Helmholtz forward and inverse problem [22, 24]. Particularly relevant to our paper is the Neumann Series Neural Operator (NSNO) approach of Chen et al. and Liu et al. [4, 17]. The NNSO is designed using U-net with the FNO or Fourier Neural Operator scheme to approximate the Born operator. This can then be used to apply the Neumann series to compute forward scattering. To solve the inverse problem, the resulting forward approximation can be combined with an optimization algorithm to solve the inverse problem.
@@
The application of the regularized inverse Born approximation using specially designed neural networks is considered in [26]. The paper also features a mathematical analysis of the generalization and approximation error.
@@
In this paper, we focus on the problem of extending the applicability of the Born approximation to cases where the weak-scattering approximation fails. Our goal is to broaden the applicability of the Born approximation and arrive at a direct quantitative estimate of the scatterer without the need to compute the solution of a nonlinear inverse problem. We investigate two approaches using a supervised convolutional neural network (CNN) scheme to correct the Born approximation. In particular, we use a generic CNN but optimize it for each case studied. The first approach–termed Born-CNN (BCNN)–performs a regularized inversion of the Born approximation applied to the far field scattering data and then uses a trained CNN to correct the resulting image. The second approach–termed CNN-Born (CNNB)–trains the neural network to pre-correct the scattering data and then applies a regularized inversion of the Born approximation to produce a corrected reconstruction. The inclusion of the Born approximation in the two approaches is motivated by the excellent fidelity of the Born approximation when weak scattering holds, so providing some physics information to the inversion. To provide a comparison, we also present results for a generic CNN applied to the inverse scattering problem along the lines of the seminal paper of [10].
@@
The novelty of this paper is to suggest and test two schemes combining the Born approximation with a neural network to extend the domain of applicability of the Born approximation. In particular, we compare the resulting predictions to the standard regularized inverse Born approximation, and to a CNN trained simply to invert the data along the lines of [10]. After optimizing the networks and training on simple data generated by a few circular scatterers, we show that the so called CNNB model is remarkably stable to added noise.
@@
Limited testing on more complex, out-of-distribution scatterers reveals that all methods improve the fidelity of the reconstructions for strong scatterers, but the inclusion of the Born approximation in the two proposed schemes produces generally better results. The study suggests that combining CNNs and the Born approximation has promise in solving inverse scattering problems and that the training phase does not need to include close copies of the scatterers.
@@
The remainder of this paper proceeds as follows. In Section 2, we briefly outline the forward and inverse problems underlying this study, summarize the Born approximation, and detail our discretization. Then, in Section 4, we give details of the architectures for the three models considered in the paper and discuss training and testing. Data generation and the main results of the paper are given in Section 5. We draw some conclusions in Section 6.
@@ 2 The Inverse scattering problem
$u^{i}$$u^{s}$$\Omega$$\eta(\textbf{x})$
Figure 1: An illustration of the scattering problem. The incident field $u^{i}$ travels with direction of propagation $\varphi$ and interacts with the compactly supported scatterer given by $\eta(\textbf{x})\not=1$ shown in blue, and contained in a bounded set $\Omega$ (shown in gray). The scattered field $u^{s}$ propagates outward with direction angle $\theta$.
We suppose that a known incident plane wave with angle of propagation $\varphi$ is given by:
$u^{i}(\mathbf{x},\mathbf{d})=\exp(\mathrm{i}\mkern 1.0muk\mathbf{x}\cdot\mathbf{d}),\quad\mathbf{d}=\langle\cos\varphi,\sin\varphi\rangle$ (1)
strikes a bounded penetrable scatterer. The square of the refractive index for the medium in which the wave propagates is denoted $\eta(\mathbf{x})\in L^{\infty}(\mathbb{R}^{2})$. This bounded function is assumed to satisfy $\Re(\eta(\mathbf{x}))\geq\eta_{\rm min}>0$, where $\eta_{\rm min}$ is a constant, and $\Im(\eta(\mathbf{x}))\geq 0$ for almost all $\textbf{x}\in\mathbb{R}^{2}$. In addition, we assume that $|\eta(\mathbf{x})|\leq\eta_{\rm max}$ a.e. for $\mathbf{x}\in\mathbb{R}^{2}$ where $\eta_{\rm max}$ is a constant. The boundedness of the scatterer implies that the contrast $\mu(\textbf{x})\mathrel{\mathop{:}}=\eta(\textbf{x})-1=0$ if $|\textbf{x}|>R$ for some $R>0$ (see, for example [6, 14]).
For a given wave number $k$, the total field $u\mathrel{\mathop{:}}=u(\mathbf{x},\mathbf{d})$ and the scattered field $u^{s}\mathrel{\mathop{:}}=u^{s}(\mathbf{x},\mathbf{d})$ satisfy the Helmholtz equation:
@@
[…]
uniformly in $\hat{\mathbf{x}}:=\mathbf{x}/|\mathbf{x}|\in\mathbb{S}^{1}$. Here, $\mathbb{S}^{1}$ denotes the unit circle in $\mathbb{R}^{2}$.
Under the conditions given above, Equations (2)-(4) have a unique solution for any $k>0$ [6]. The Forward Problem consists of solving the above linear well-posed problem given $\mathbf{d}$, $k$, and $\eta$ (see Fig. 1).
It follows from the fact that $u^{s}$ satisfies the Helmholtz equation and the radiation condition that $u^{s}$ exhibits an asymptotic expansion as an outgoing cylindrical wave for $|\mathbf{x}|$ sufficiently large:
$u^{s}(\textbf{x},\textbf{d})=\frac{\exp(\mathrm{i}\mkern 1.0muk|\textbf{x}|)}{\sqrt{|\textbf{x}|}}\left(u_{\infty}(\hat{\textbf{x}},\textbf{d})+\mathcal{O}(|\textbf{x}|^{-1})\right)\text{ as }|\textbf{x}|\to\infty,$ (5)
where $u_{\infty}:\mathbb{S}^{1}\to\mathbb{C}$ is called the far field pattern of the scattered wave [6].
@@
In this study, we use synthetic scattering data generated through a standard finite element approach to approximating (2)-(4). For each incident direction d, the total field $u$ in a neighborhood of the scatterer is computed using the Netgen package [21] with 4th-order elements and a mesh-size request of one-eighth of the local wavelength of the wave. The boundary of each scatterer is fitted using isoparametric curved elements. The Sommerfeld radiation condition is handled through a radial Perfectly Matched Layer (PML) implemented by Netgen using complex stretching as discussed in [5]. To generate an approximate far field pattern, we follow [18] to map the near field to a far field pattern.
@@ 2.1 The Born approximation
$H^{1}_{loc}(\mathbb{R}^{2})=\{u:\mathbb{R}^{2}\to\mathbb{C}\>|\>u|_{B_{t}}\in H^{1}(B_{t})\>\>\forall\>\>t>0\},$ (6)
where $B_{t}=B(t,0)$ denotes a disk of radius $t$ centered at the origin. In particular, let $\Omega=B_{R}$ be the smallest such disk containing the support of $\mu$. It can be shown that if $u\in H^{1}_{loc}(\mathbb{R}^{2})$ is a solution to the scattering problem (2)-(4), then $u|_{\Omega}\in L^{2}(\Omega)$ and satisfies the Lippmann-Schwinger equation:
$u(\textbf{x})=u^{i}(\textbf{x})+k^{2}\int_{\Omega}\Phi(\textbf{x},\textbf{y})\mu(\textbf{y})u(\textbf{y})\,d\textbf{y}\mathrel{\mathop{:}}=u^{i}(\textbf{x})-(Tu)(\textbf{x})\quad\forall\>\>\textbf{x}\in\mathbb{R}^{2},$ (7)
@@
$u(\textbf{x})=\sum_{j=0}^{\infty}(-1)^{j}T^{j}u^{i}$ (8)
provided that the operator norm $\|T\|_{\infty}<1$. A sufficient condition for this is [19]
@@
where $\|\cdot\|_{\infty}$ is the max norm and
@@
When $\|T\|_{\infty}<1$, the first two terms of (8) provide the Born approximation of the field $u$:

### v3

@@ Abstract
Time-harmonic acoustic inverse scattering concerns the ill-posed and nonlinear problem of determining the refractive index of an inaccessible, penetrable scatterer based on far field wave scattering data. When the scattering is weak, the regularized inverse Born approximation provides a linearized model for recovering the shape and material properties of a scatterer. We propose two convolutional neural network (CNN) algorithms to correct the traditional inverse Born approximation even when the scattering is not weak. These are denoted Born-CNN (BCNN) and CNN-Born (CNNB). BCNN applies a post-correction to the Born reconstruction, while CNNB pre-corrects the data. Both methods leverage the Born approximation’s excellent fidelity in weak scattering, while extending its applicability beyond its theoretical limits. CNNB particularly exhibits a strong generalization to more complex out of distribution scatterers. Based on numerical tests and benchmarking against other standard approaches, our corrected Born models provide alternative data-driven methods for obtaining the refractive index, extending the utility of the Born approximation to regimes where the traditional method fails.
00footnotetext: Funding: The research of A.D. is supported by the University of Delaware Undergraduate Research Program. The research of T.L. is partially supported by the Research Council of Finland via the Finnish Center of Excellence of Inverse Modeling and Imaging, the research project 321761, and the Flagship of Advanced Mathematics for Sensing Imaging and Modeling grant 358944. The research of P.M. was partially supported by the US AFOSR under grant number FA9550-23-1-0256.
@@ 1 Introduction
We shall study a particular inverse scattering problem: the inverse medium problem for the Helmholtz equation. In this case, we seek to reconstruct the refractive index of a bounded penetrable scatterer from far field acoustic data. The major complication is that this inverse scattering problem is both ill-posed and nonlinear [10, Theorem 4.21, page 448]. Current approaches can be divided into two broad classes: 1) quantitative methods and 2) qualitative methods. A quantitative method attempts to reconstruct the scatterer directly. Often this involves a non-linear and regularized optimization problem that seeks a reconstruction corresponding to a far field pattern that matches measurement data. This is computationally intensive and can fail due to local minima [10]. Another quantitative approach is to assume that the scattering is weak and to use the inverse Born approximation. We shall discuss this in more detail shortly since correcting this approach is the subject of our paper. In contrast to these quantitative methods, the goal of a qualitative method such as the Linear Sampling Method [5, 10] is to approximate the support or boundary of the scatterer. Approaches of this type do not involve optimization or the solution of the forward problem. They do not require strong a priori assumptions, but cannot directly distinguish different materials in the field of view.
@@
Turning now to the Born approximation, the forward scattering problem that maps a known refractive index to the predicted far field pattern is given by a Neumann series in a certain integral operator, provided an appropriate norm of the integral operator is less than one. Conditions for this to occur have been derived in several cases (for example, [18] in the seismic context) and we recall a simple sufficient condition in Section 2. Selecting only the first term in this series defines a linear map, the Born approximation, from the contrast to the approximate far field pattern.
@@
For the inverse problem, the inverse Born approximation can be understood as inverting the linearized Born approximation. Inverting this linear map is ill-posed, but removes the difficulty of nonlinearity and provides an avenue to solve the inverse problem approximately using regularization to restore stability. An early reference for this technique is  [4], while applications and computational techniques are described in [12, 13]. The inverse Born approximation has been widely applied to neutron scattering, medical imaging, and seismic inverse problems [18]. Within the weak scattering approximation, there has been a great deal of work to incorporate higher-order expansions into the forward and inverse Born approximation (see the review of Moskow and Schotland [34]). Our goal in this paper is to correct the Born approximation and extend its applicability beyond the weak scattering limit using neural networks (NNs) as correctors.
While traditional inverse scattering techniques offer interpretability and theoretical grounding, they often suffer from high computational cost or restrictive assumptions (e.g., weak scattering). To address these limitations, recent years have seen rapid growth in the use of machine learning, particularly neural networks, to learn the inverse map directly or to enhance physics-based approximations.
@@
One machine learning approach to correcting the Born approximation is the statistical approximation error correction used in [20], which shows that the Born approximation can be extended outside the weak scattering approximation by a suitable training approach if the scatterer is well represented in the training data. The approach of [20] is based on Bayesian statistics and not on neural networks.
We now turn to a review of recent efforts to use NN based machine learning techniques to address the inverse scattering problem. These approaches vary in architecture and in whether they incorporate aspects of the underlying physics of wave propagation. We emphasize that this is not intended to be a comprehensive review of this rapidly evolving field, but rather a focused overview of representative approaches most relevant to our work.
Viewing the map from the far field to the regularized reconstruction of the scatterer as an unknown nonlinear function, we can appeal to the universal approximation property (first proved in [11]) of NNs to approximate this map. Note that regularization is required because, without it, the map is not defined [10].
For the inverse problem, we can distinguish two general approaches: 1) Use a NN to approximate the forward problem and then use this model in a traditional optimization framework for the inverse problem (see for example [39, 36, 6, 28]), 2) Develop end-to-end algorithms that seek to solve the inverse problem directly by approximating the above mentioned data to coefficient map.
We are interested in the second approach. Within the end-to-end paradigm, an important approach is based on Physics Informed Neural Networks (PINNs) [36, 8, 7, 35]. Here the governing differential equation is incorporated in the training. Although it is possible to solve infinite domain scattering problems using PINNs (for example [42]), we are not aware of a PINNs approach to solving the inverse scattering problem using far field data. Instead, we shall use the inverse Born approximation to supply physics information to the networks we design.
Apart from PINNs, there is an extensive literature on end-to-end inversion using supervised NNs including [21, 43, 15, 2, 31, 46, 44]. In particular, this is the approach in [15] which is inspired by the Born approximation and is a motivating paper for our work. They use their BCR-Net [14] to approximate the inverse problem since it is based on a nonstandard form of a wavelet representation described in [3] and designed to approximate nonlinear integral operators. We shall compare our models to the BCR-Net approach in Section 5.
@@
Another end-to-end network motivated by the regularized inverse Born approximation for weak scattering data is considered in [46]. That paper also features a mathematical analysis of the generalization and approximation error, and numerical results showing good approximation on weak scattering data.
More generally, beyond architectures motivated by a specific physics-based representation, there has been increasing interest in neural operator frameworks such as the Fourier Neural Operator (FNO) and Deep Operator Network (DeepONet). These approaches learn mappings between infinite-dimensional function spaces, enabling mesh-independent generalization and rapid solution of PDE-based forward and inverse problems [24, 25, 26, 30, 41]. Building on these ideas, a recent hybrid framework termed the Neural Inverse Operator (NIO) combining FNOs and DeepONets was introduced by Molinaro et al. [31]. This has been applied to a wide range of inverse coefficient identification problems, showing excellent results. However, the inverse scattering problem considered in that paper uses a bounded domain and Dirichlet and Neumann data pairs, so it is not directly comparable to our upcoming model using far field data. Nonetheless, with small changes, we can train NIO on far field data and compare results with our Born based schemes. Note that these methods may suffer from hallucinations [17].
Our approach is inspired by the strengths and limitations of the Born approximation, which provides valuable physics-informed structure, especially in the weak scattering regime, but fails for stronger contrasts. Rather than discarding it entirely, we propose to integrate it into supervised learning pipelines as an inductive bias, thereby combining the interpretability of physics-based models with the flexibility of neural networks. It is also hoped that this approach will reduce the tendency to hallucinate noted in [17].
@@
Our goal is to broaden the applicability of the Born approximation and arrive at a direct quantitative estimate of the scatterer without the need to compute the solution of a nonlinear inverse problem. We investigate two approaches using a supervised convolutional neural network (CNN) scheme to correct the Born approximation. In particular, we use a generic CNN, but optimize it for each case studied. The first approach–termed Born-CNN (BCNN)–performs a regularized inversion of the Born approximation applied to the far field scattering data and then uses a trained CNN to correct the resulting image. The second approach–termed CNN-Born (CNNB)–trains the neural network to pre-correct the scattering data and then applies a regularized inversion of the Born approximation to produce a corrected reconstruction.
@@
The novelty of this paper is to suggest and test the above mentioned schemes combining the Born approximation with a CNN to extend the domain of applicability of the Born approximation. In particular, we compare the resulting predictions to those from the standard regularized inverse Born approximation, to a slightly modified version of the Neural Inverse Operator (NIO) from [31], to the BCR-Net network described in [14], and to a basic CNN. After optimizing the networks and training on simple data generated by a few circular scatterers, we test the networks on several out of distribution scenarios.
@@
The study suggests that combining CNNs and the Born approximation has promise in solving inverse scattering problems and that the training phase does not need to include close copies of the scatterers.
@@
The remainder of this paper proceeds as follows. In Section 2, we briefly outline the forward and inverse problems underlying this study, summarize the Born approximation, and detail our discretization. Then, in Section 4, we give details of the architectures for the three models considered in the paper and discuss training and testing. Data generation and the main numerical results of the paper are given in Section 5. We draw some conclusions in Section 6.
@@ 2 The Inverse scattering problem
We suppose that a known incident plane wave with angle of propagation $\varphi$ is given by
$u^{i}(\mathbf{x};\mathbf{d})=\exp(\mathrm{i}\mkern 1.0muk\mathbf{x}\cdot\mathbf{d}),\quad\mathbf{d}=\langle\cos\varphi,\sin\varphi\rangle$ (1)
and that it strikes a bounded penetrable scatterer.
The square of the refractive index for the medium in which the wave propagates is denoted $\eta(\mathbf{x})\in L^{\infty}(\mathbb{R}^{2})$. This bounded function is assumed to satisfy $\Re(\eta(\mathbf{x}))\geq\eta_{\rm min}>0$, where $\eta_{\rm min}$ is a constant, and $\Im(\eta(\mathbf{x}))\geq 0$ for almost all $\textbf{x}\in\mathbb{R}^{2}$. In addition, we assume that $|\eta(\mathbf{x})|\leq\eta_{\rm max}$ a.e. for $\mathbf{x}\in\mathbb{R}^{2}$ where $\eta_{\rm max}$ is a constant. The boundedness of the scatterer implies that the contrast $\mu(\textbf{x})\mathrel{\mathop{:}}=\eta(\textbf{x})-1=0$ if $|\textbf{x}|>R$ for some $R>0$ (see, for example [10, 22]).
For a given wave number $k$, the total field $u\mathrel{\mathop{:}}=u(\mathbf{x};\mathbf{d})$ and the scattered field $u^{s}\mathrel{\mathop{:}}=u^{s}(\mathbf{x};\mathbf{d})$ satisfy the Helmholtz equation:
@@
$r^{1/2}\left(\frac{\partial u^{s}}{\partial r}-\mathrm{i}\mkern 1.0muku^{s}\right)$ $\to 0\text{ as }r\mathrel{\mathop{:}}=|\mathbf{x}|\to\infty,$ (4)
uniformly in $\hat{\mathbf{x}}:=\mathbf{x}/|\mathbf{x}|\in\mathbb{S}^{1}$ where $\mathbb{S}^{1}$ denotes the unit circle in $\mathbb{R}^{2}$.
$u^{i}$$u^{s}$$\Omega$$\eta(\textbf{x})$
Figure 1: An illustration of the scattering problem. The incident field $u^{i}$ travels with direction of propagation $\mathbf{d}$ and interacts with the compactly supported scatterer given by $\eta(\textbf{x})\not=1$ shown in blue, and contained in a bounded set $\Omega$ (shown in gray). The scattered field $u^{s}$ propagates outward with direction angle $\theta$.
Under the conditions given above, Equations (2)-(4) have a unique solution for any $k>0$ [10]. The Forward Problem consists of solving the above linear well-posed problem given $\mathbf{d}$, $k$, and $\eta$ (see Fig. 1). It follows from the fact that $u^{s}$ satisfies the Helmholtz equation and the radiation condition that $u^{s}$ exhibits an asymptotic expansion as an outgoing cylindrical wave for $|\mathbf{x}|$ sufficiently large:
$u^{s}(\textbf{x};\textbf{d})=\frac{\exp(\mathrm{i}\mkern 1.0muk|\textbf{x}|)}{\sqrt{|\textbf{x}|}}\left(u_{\infty}(\hat{\textbf{x}},\textbf{d})+\mathcal{O}(|\textbf{x}|^{-1})\right)\text{ as }|\textbf{x}|\to\infty,$ (5)
where $u_{\infty}:\mathbb{S}^{1}\times\mathbb{S}^{1}\to\mathbb{C}$ is called the far field pattern of the scattered wave [10].
@@
In this study, we use synthetic scattering data generated through a standard finite element approach to approximating (2)-(4). For each incident direction d, the total field $u$ in a neighborhood of the scatterer is computed using the Netgen package [37] with 4th-order finite elements and a mesh-size request of one-eighth of the local wavelength of the wave. The boundary of each scatterer is fitted using isoparametric curved elements. The Sommerfeld radiation condition is handled through a radial Perfectly Matched Layer (PML) implemented by Netgen using complex stretching as discussed in [9]. To generate an approximate far field pattern, we follow [32] to map the near field to a far field pattern. The implementation of our forward solver can be found at https://github.com/nibj/Helmholtz-scattering-data.
@@ 2.1 The Born approximation
$H^{1}_{loc}(\mathbb{R}^{2})=\left\{u:\mathbb{R}^{2}\to\mathbb{C}\>|\>u|_{K}\in H^{1}(K)\>\forall\mbox{ compact subsets }K\subset\mathbb{R}^{2}\right\}.$ (6)
In particular, let $\Omega$ denote a compact domain containing the support of $\mu$ (the sets where $\mu\not=0$). It can be shown that if $u\in H^{1}_{loc}(\mathbb{R}^{2})$ is a solution to the scattering problem (2)-(4), then $u|_{\Omega}\in L^{2}(\Omega)$ and satisfies the Lippmann-Schwinger equation [10]:
$u(\textbf{x};\textbf{d})=u^{i}(\textbf{x};\textbf{d})+k^{2}\int_{\Omega}\Phi(\textbf{x},\textbf{y})\mu(\textbf{y})u(\textbf{y};\textbf{d})\,d\textbf{y}\mathrel{\mathop{:}}=u^{i}(\textbf{x};\textbf{d})-(Tu)(\textbf{x};\textbf{d})\quad\forall\>\textbf{x}\in\mathbb{R}^{2},\forall\>\textbf{d}\in\mathbb{S}^{1}$ (7)
@@
$u(\textbf{x};\textbf{d})=\sum_{j=0}^{\infty}(-1)^{j}(T^{j}u^{i})(\textbf{x};\textbf{d})$ (8)
provided that the operator norm $\|T\|_{2}<1$. A sufficient condition for this is [33]
@@
where $\|\cdot\|_{\infty}$ is the max norm and $B_{R}$ is the smallest ball, with radius $R$, containing the support of the scatterers and
@@
When $\|T\|_{2}<1$, the first two terms of (8) provide the Born approximation of the field $u$: